OpenStudy (naudiapie):

Sam Collins wants to get the best deal in cereal he can find. He finds rice puffs for $2.49 for 24 ounces, $2.69 for 36 ounces, $2.29 for 12 ounces, and $4.00 for 48 ounces. What is the unit price of the best buy?

OpenStudy (johnweldon1993):

You know the rules! Tell me! :P

OpenStudy (naudiapie):

but I'm lazy

OpenStudy (naudiapie):

jk

OpenStudy (johnweldon1993):

But but but...you get presents after :O

OpenStudy (naudiapie):

OMG BALLOONS

OpenStudy (naudiapie):

I'm probably going to need you to help me every day D:

OpenStudy (johnweldon1993):

Well it's a good thing I found you then :O <3

OpenStudy (naudiapie):

am i just dividing

OpenStudy (johnweldon1993):

mmhmm just like last time...list each result you get :)

OpenStudy (naudiapie):

well (2.49)/(24) = 0.10

OpenStudy (naudiapie):

do that for the rest?

OpenStudy (johnweldon1993):

mmhmm ^_^

OpenStudy (naudiapie):

but look at the answers

OpenStudy (naudiapie):

Sam Collins wants to get the best deal in cereal he can find. He finds rice puffs for $2.49 for 24 ounces, $2.69 for 36 ounces, $2.29 for 12 ounces, and $4.00 for 48 ounces. What is the unit price of the best buy? 7.47¢ per ounce 8.37¢ per ounce 8.33¢ per ounce 9.08¢ per ounce

OpenStudy (johnweldon1993):

mmhmm you'll see it when you list em :D

OpenStudy (naudiapie):

2.69/36 = 0.074 2.29/12= 0.19 4/48 = 0.08 2.49/24 = 0.10

OpenStudy (johnweldon1993):

Perfect! so which one is the smallest???? the 0.074 one right? Remember that is Dollars...so in cents...that would be times 100 or 7.4 cents

OpenStudy (naudiapie):

wow okay i see now, yeah i guessed on that answer twice and didn't choose 7.4

OpenStudy (johnweldon1993):

No guessing when I'm here >.< :P

OpenStudy (naudiapie):

pwease?

OpenStudy (johnweldon1993):

No >.< No you only get 1 balloon!!!

OpenStudy (naudiapie):

/cries

OpenStudy (naudiapie):

well give me it then

OpenStudy (johnweldon1993):



OpenStudy (naudiapie):